# DLinear-based Prediction of Remaining Useful Life of Lithium-Ion Batteries: Feature Engineering through Explainable Artificial Intelligence

### v1

@@ 3.2.2 SHAP-based approach
Where $\phi_{i}$ is the Shapley value of the $i$-th feature Fi, $N$ is the set of all features (e.g., in our case, $\left|N\right|$ is 20, see the Table 1), $N\setminus\left\{i\right\}$ is the set of all possible combinations of remaining features excluding feature Fi, $S$ is a subset of $N\setminus\left\{i\right\}$, $f\left(S\cup\left\{i\right\}\right)$ is the model prediction of the union of Fi and feature set $S$, and $f$($S$) is the model prediction for the features in $S$.
@@ 4 Results and discussion
Figure 3: PCC of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
Figure 4: SHAP score of all parameters and RUL for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
The current, voltage, temperature, and time data collected from the cycling experiment were reproduced into 20 features as shown in Table 1. Fig. 3 presents the results of PCC on 20 features for four cells. The three most sensitive features for each cell, as identified by the PCC, are:
@@
B0006, B0007, and B0018 include voltage characteristics during the discharging stage and time-related parameters during the charging or discharging stages as key parameters. In contrast, all three key parameters for B0005 are related to temperature characteristics, demonstrating significant variability across the four cells.
Fig. 4 shows the results of SHAP analysis, that F2 (Variance of measured voltage in discharging) and F11 (Variance of temperature in discharging) were selected as the highest values across all four cells. In addition to F2 and F11, F10 (Average temperature in discharging) was identified as one of the three most sensitive feature for B0005. Meanwhile, F18 (CC charging time), along with F2 and F11, was identified as one of the three most sensitive feature for B0006, B0007, and B0018.
When training with the existing 70 cycles, the EoL prediction gap was well captured in Fig. 5. EoL was defined as 80% of the nominal capacity. In the case of LSTM, it captured and tracked the decreasing trend well after the training cycles. However, as the prediction progressed toward the later stages, it failed to follow the downward trend, resulting in a flat graph. This is because of the gradient vanishing problem mentioned previously.
The performance of Transformer was found to be inferior than LSTM due to the issue of over-stationarization, which is considered a drawback of the Transformer algorithm [38]. Since the Transformer was originally developed for natural language processing rather than time-series data, it fails to capture critical temporal dependencies in non-stationary time-series data with clear downward trends. Despite the fact that forecasting non-stationary time-series data is crucial for real-world applications including battery RUL predictions, such over-stationarization limits the prediction capability of the Transformer in these non-stationary time-series data. Previous studies on RUL prediction using Transformers [23, 39] have demonstrated that Transformers outperform LSTM in terms of predictive performance with revised model of Transformer. However, this study focuses on a comprehensive analysis of feature engineering considering CtCV and the improvement in RUL prediction performance driven by the characteristics of the DLinear algorithm. In this context, optimization of the model architecture was not considered in this study.
While LSTM and Transformer show a significant decrease in prediction performance as the training period decreases, DLinear maintained relatively accurate EoL prediction performance. Even when the training cycles were reduced to a maximum of 30 cycles, The performance of DLinear was observed to decline. Thus, DLinear demonstrated excellent performance in battery RUL prediction by separating the trend and seasonality of time-series data and learning them linearly, thereby effectively overcoming both the gradient vanishing problem of LSTM and the over-stationarization problem of the Transformer.
Figure 5: RUL prediction of four LIBs : (a) B0005, (b) B0006, (c) B0007, (d): B0018
@@
Figure 6: Comparison of R2 score according to algorithm and training cycle
$R^{2}=1-\frac{SS_{\text{res}}}{SS_{\text{tot}}}$ (5)
Fig. 6 illustrates the R2 scores of the models based on each algorithm, while Table 2 presents the MSE values according to the algorithm and the size of the training cycles, and Table 3 shows the corresponding R2 scores. The R2 score is not affected by the scale of the data, therefore particularly useful for evaluating battery data with varying training cycle lengths and total cycle durations. In contrast, MSE, by squaring the errors, sensitively captures large errors in time-series data and is differentiable, making it a commonly used loss function in machine learning algorithms. When evaluated using MSE and R2 score, DLinear’s performance was outstanding. The bolded values in Table 2 and Table 3 indicate the superior RUL prediction performance of DLinear. Even when the training cycles were reduced to 40, DLinear achieved an R2 score of 0.995. $SS_{\text{res}}$ is the sum of squared differences between the observed values and the predicted values, and $SS_{\text{tot}}$ is the sum of squared differences between the observed values and the mean of the observed data. In cases where the model performs worse than simply predicting the mean of the observed values, the R2 score will be negative. This is because Eq. 5 subtracts the ratio of the residual sum of squares from 1, meaning the resulting negative value is not constrained within the range of -1 to 1.
Figure 7: RUL prediction of B0005 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 8: RUL prediction of B0006 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 9: RUL prediction of B0007 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Figure 10: RUL prediction of B0018 according to training cycle : (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d): 30 cycle
Table 2: Comparison of MSE by algorithm according to training cycle
Algorithm Training Cycle
70 50 40 30
B0005 LSTM 0.03407 0.01821 0.01738 0.01999
Transformer 0.15731 0.31750 0.11027 0.37554
DLinear 0.00003 0.00009 0.00019 0.00042
B0006 LSTM 0.02038 0.00924 0.01089 0.01532
Transformer 0.03699 0.24336 0.60980 0.35563
DLinear 0.00003 0.00006 0.00009 0.00019
B0007 LSTM 0.04070 0.01757 0.01530 0.01355
Transformer 0.02553 0.19942 0.10073 0.15515
DLinear 0.00007 0.00031 0.00062 0.00220
B0018 LSTM 0.01101 0.03776 0.02583 0.01577
Transformer 0.02165 0.02092 0.07549 0.20434
DLinear 0.00002 0.00003 0.00003 0.00006
Table 3: Comparison of R2 by algorithm according to training cycle
Algorithm Training Cycle
70 50 40 30
B0005 LSTM -0.17398 0.66193 0.74823 0.75999
Transformer -2.10452 -3.59821 -0.59773 -1.68900
DLinear 0.98387 0.99705 0.99628 0.68536
B0006 LSTM -0.37580 0.63463 0.69113 0.66219
Transformer -1.49730 8.62568 16.29935 6.84430
DLinear 0.99772 0.34021 -0.73222 -0.96263
B0007 LSTM -0.69123 0.61148 0.74657 0.81998
Transformer -0.06079 3.40917 -0.66853 1.06123
DLinear 0.97061 0.99422 0.99511 0.33864
B0018 LSTM -0.57772 -0.40026 0.38087 0.67313
Transformer -2.10273 0.22419 -0.80949 3.21659
DLinear 0.99418 0.98418 0.99464 0.95958
In PHM technologies, the ultimate goal is to achieve accurate predictions with fewer training cycles as possible. To evaluate whether this study can contribute to PHM technology, we progressively reduced the training cycles to 70, 50, 40, and 30 and analyzed the prediction performance. Fig. 7, 8, 9 and, 10 present the prediction results for four different battery cells (B0005, B0006, B0007, and B0018, respectively) while the number of training cycles was reduced. The corresponding MSE and R2 values are shown in Table 2 and Table 3. Reducing training cycles, LSTM exhibited poor performance, with its predictions deviating significantly from the actual values in the later stages, as shown in Fig. 5. Similarly, due to the over-stationarization issue, predictions made using the Transformer also demonstrated severe fluctuations which was to minimize the loss function, consistent with the behavior observed in Fig. 5. On the other hand, except for B0006 cell (Fig. 8), DLinear demonstrates superior prediction performance compared to LSTM and Transformer even at 50 cycles and 40 cycles. However, except for B0018 (Fig. 10), there is a considerable difference between the real cycling behavior and the predicted cycling behavior when the training cycles are set to 30. In fact, recent research using Transformer demonstrates superior predictive performance compared to LSTM [40, 23]. However, this study focuses on feature engineering, and by using a simple network without considering improvements to the machine learning algorithms or hyperparameter tuning, it makes it easier to reproduce the results or apply this framework to other battery systems.
SHAP and PCC were employed for feature selection. Although using the three most sensitive features of performance of PCC resulted in higher prediction performance, there was a lack of consistency across different batteries (B0005, B0006, B0007, B0018, respectively). In contrast, while the three most sensitive features identified using SHAP yielded slightly lower prediction performance than the performance of PCC, they demonstrated consistency in feature selection across different batteries. This result suggests that SHAP-based feature selection has a higher potential for transfer learning between batteries, suggesting a certain level of robustness. In battery capacity research, sampling is often conducted to ensure representativeness that accounts for differences between cells, underscoring the importance of generalization performance. The consistent selection of the same features across multiple batteries indicates that these features are likely useful in a general context rather than being specific to individual cells. This consistency also implies potential reductions in future data collection and analysis costs, making it significant for broader applications.
Figure 11: RUL prediction of B0005 cell according to feature selection: (a) PCC, (b) SHAP
Table 4: The most sensitive and least sensitive features in terms of importance, as determined by SHAP analysis
B0005 B0006 B0007 B0018
Three most sensitive 2a,*, 10, 11a 2a, 11a, 18a 2a, 11a, 18a 2a, 11a, 18a
Three least sensitive 1, 7, 20 1, 5b, 14b 5b, 14b, 16 5b, 14b, 16
* Based on the feature selection results using SHAP, the extracted features, denoted by superscript a, are identified as key features in three or more cells, while those, denoted by superscript b, are identified as not significant in three or more cells.
Table 5: The three most sensitive and three least sensitive features in terms of importance, as determined by PCC
B0005 B0006 B0007 B0018
Three most sensitive 10, 13, 12 18, 20, 3 7, 8, 18 7, 8, 20
Three least sensitive 17, 19, 20 13, 17, 15 13, 14, 15 10, 14, 16
Fig. 11 show the prediction results for B0005 cell using the sets of the three most sensitive features and the three least sensitive features, as analyzed through PCC and SHAP, respectively. The sets of the three most sensitive parameters and the three least sensitive parameters, as determined by SHAP and PCC analyses, are presented in Table 4 and Table 5.
SHAP analysis reveals that more than half of the three most sensitive features are shared among batteries. A noteworthy point is, the key features identified by PCC differ significantly across the four cells, whereas those identified by SHAP analysis are identical for B0006, B0007, and B0018, wich B0005 sharing two features with the other cells. A possible interpretation of these results is that the key parameters identified through SHAP-based sensitivity analysis provide robust results concerning CtCV. In practice, the uncertainties propagated by various causes, such as differences in electrode thickness during manufacturing stages (e.g., Calendaring process), imply that cycling behavior can vary even when cycling the same battery system under identical conditions [5]. The high robustness of SHAP suggests that it is a promising approach for predicting the RUL of large-capacity batteries (e.g., EVs) used in modules or packs.
Meanwhile, as shown in Table 5, PCC revealed cases where a feature ranked low for one battery was ranked high for another. For example, F10, F13, and F20 were selected as the three most sensitive features based on PCC for some batteries, while they were in the three least sensitive features for others. These three features are related to temperature and time, and since batteries are significantly influenced by their surroundings, especially the ambient temperature, measurements such as maximum or minimum values can be affected by external factors or measurement errors. Therefore, these features appear to lack consistency in their significance across different batteries when evaluated based on PCC.
Figure 12: RUL prediction of B0005 cell using two most sensitive features identified by SHAP: F2, F11
In Table 4, F2 and F11 were consistently selected as the three most sensitive features across all four batteries. Building on this observation, we also examined the performance of predictions using only these two features. However, the performance did not match that of predictions made using three features. This suggests that, in addition to F2 and F11, including additional features such as F18, which encapsulate lifespan-related information, is recommended for improved prediction accuracy.
@@ 5 Conclusion
PHM technology involves real-time diagnosis and prediction of the health state of a target system. Accurate RUL prediction helps implement condition-based management, where interventions are taken only when necessary by predicting the preemptive system replacement point before failures or defects occur. However, predicting failures through RUL requires measuring charge/discharge data during real usage, which is impractical and time-consuming during experiments. Therefore, it’s crucial to predict with higher accuracy using fewer experiments. In cases of clear trend-based predictions like battery degradation, DLinear outperforms models utilizing other algorithms. Even when trained on the same number of cycles, DLinear can predict further and more accurately than conventional LSTM or Transformer models. This suggests its potential contribution to implementing Condition Based Management for battery PHM technology in real industrial applications. On the other hand, identifying key features for predicting RUL aids in the implementation of PHM technology. PCC yields good performance for individual cells, but the key features vary significantly due to CtCV. Additionally, SHAP does not perform as well as PCC in terms of overall performance. However, SHAP consistently identifies key features, making it advantageous for battery research requiring generalization, despite the CtCV. Particularly, CtCV arises from both intra-cell factors such as morphology, architecture, and composition, and inter-cell factors like winding, connection, and assembly processes. Therefore, creating homogeneous batteries is challenging. In this regard, it is noteworthy that SHAP evaluates the importance of the same features despite CtCV.

### v2

@@ 4 Results and discussion
4.1 SOH prediction
This work aims to provide insight into the SOH prediction research field by comparing widely used methods for processing data collected through BMS and generating prediction models for SOH prediction. In particular, many works are using PCC for feature engineering [45, 46, 47]. In fact, we believe that global sensitivity analysis methods, such as SHAP and Sobol’ indices, can comprehensively analyze the influence of parameters, just like PCC. In addition to widely used time-series prediction models such as RNN, LSTM, and Transformer, advanced machine learning models were generated by integrating time-series prediction models with various machine learning algorithms, such as Convolutional Neural Network (CNN), and applied to SOH prediction (Table 2). Here, the DLinear, which has strengths in learning patterns of time-series data, is applied, and its accuracy is compared with that of LSTM and Transformer. According to Zeng et al. [31], the performance of the DLinear is far superior to that of the Transformer model in capturing temporal dynamic patterns between the set of continuous points. The SOH of a battery often shows a clear decreasing trend even when considering the capacity increase phenomenon in early cycles. Therefore, we believe that DLinear may be a suitable method for SOH prediction. However, for the fundamental performance comparison of the three algorithms, advanced models through integration with other ML algorithms are not considered.
Table 2: Literature on advanced ML models for SOH prediction
Reference Methodology Hyperparameter optimization
Fu et al. [48] Variational mode decomposition(VMD)-Permutation entropy(PE)-Improved dung beetle optimization (IDBO)-Temporal convolutional network(TCN) IDBO
Xia et al. [49] Bi-directional gated recurrent unit (BiGRU) Particle swarm optimization (PSO)
He et al. [50] Gaussian process regression (GPR) Improved PSO with mutation factor and self-adaptive weight adjustment
Zhou et al. [51] Convolutional bi-directional LSTM (CNN-Bi-LSTM) neural network Sparrow search algorithm (SSA)
Zhang et al. [52] Mixed Kernel function relevance vector regression (MKRVR) Differential evolution (DE) with gray wolf optimizer (GWO)
Zheng et al. [53] GRU-TCN-Support vector machine (SVR) (GRU-TCN: Meta learner, SVR: Base learner) Whale optimization algorithm (WOA)
Tang et al. [54] CNN-BiGRU-Attention mechanism (AM) Kepler optimization algorithm (KOA)
Ma et al. [55] LSTM Differential evolution GWO (DEGWO)
Zhu et al. [56] BiLSTM
Figure 3: SOH prediction results for four cells using all the features specified in the Table 1: (a) B0005, (b) B0006, (c) B0007, (d) B0018
Fig. 3 shows the prediction results of SOH for four cells using all the features specified in Table 1. In order to compare the SOH prediction accuracy for the four cells under the same criteria, we used up to the 70th cycle as training data. In addition, SOH of the four cells is defined as the initial capacity of each cell, not the nominal capacity. In this work, the EoL is defined as reaching 80% of the initial capacity, and the initial capacity of the four cells and EoL are different due to CtCV. DLinear shows considerable prediction results for all four cells, but LSTM and Transformer do not. In particular, DLinear shows lower root mean square error (RMSE) with fewer training cycles than previous works (Table 3). Despite the considerable prediction performance of the DLinear model, the prediction using all 20 features can often cause overfitting or increase the training time, which can be the main cause.
Table 3: Comparison of RMSE across NASA battery cells with previous works
Battery Ref. [25] Ref. [57] Our work
Training cycles RMSE Training cycles RMSE Training cycles RMSE
B0005 110 0.0168 80 0.0147 70 0.0055
B0006 90 0.0152 80 0.0196 70 0.0055
B0007 140 0.0085 80 0.0110 70 0.0084
4.2 Feature engineering for SOH prediction
Fig. 4 presents the results of PCC on 20 features for four cells. The three most sensitive features for each cell, as identified by the PCC, are:
@@ 4.2 Feature engineering for SOH prediction
In the B0005 cell, the three features (F10, F12, F13) with large values of PCC are all temperature-related features. On the other hand, the remaining three cells include voltage-related features in the discharge phase and time-related features in the charge or discharge phase as key features. The reason why the correlations with SOH are different, even though all four cells are cycled by the same charge and discharge protocol, can be explained by CtCV. In particular, it is noteworthy that no feature is identified as a key feature for all four cells.
Figure 4: PCC of all features and SOH for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
Figure 5: SHAP score of all parameters and SOH for four battery cells: (a) B0005 cell, (b) B0006 cell, (c) B0007 cell, (d) B0018 cell
Fig. 5 shows the results of global sensitivity analysis based on SHAP for 20 features. The three most sensitive features for each cell, as identified by the SHAP, are:
  • B0005: F2 (Variance of measured voltage in discharging), F10 (Average temperature in discharging), F11 (Variance of temperature in discharging)
  • B0006, B0007, and B0018: F2 (Variance of measured voltage in discharging), F11 (Variance of temperature in discharging), F18 (CC charging time)
It is noteworthy that B0006, B0007, and B0018 cells are all explained by the same features in the SHAP results. In addition, B0005 also includes F2 and F11 among the three features as key features. Unlike PCC, the results in SHAP show that there is a common feature subset across the four cells, which means that SHAP can provide feature selection results that successfully reflect CtCV.
Figure 6: Comparison of results by time-series prediction model according to feature selection method and number of key features: (a) PCC-LSTM, (b) SHAP-LSTM, (c) PCC-Transformer, (d) SHAP-Transformer, (e) PCC-DLinear, and (f) SHAP-DLinear.
Fig. 6 shows the results applied to B0005 cell using 70 cycles for training to compare the accuracy of combinations of two feature selection methods and three time-series prediction models. We also investigated the effect of the number of features by training the model with three, four, and five features. This analysis aimed to validate our hypothesis that simply increasing the feature count does not guarantee better performance. For example, Fig. 6(a) and Fig. 6(d) present the results of training LSTM using features extracted from PCC and SHAP, respectively.
Firstly, for LSTM, the results are almost similar between PCC and SHAP. Also, increasing the number of features does not necessarily increase prediction accuracy. A common phenomenon in all LSTM applications is that, after several cycles in the prediction phase, a flat SOH profile emerges, indicating that the LSTM model may encounter gradient vanishing issues.
Secondly, for the Transformer model, PCC shows significantly improved predictive performance compared to LSTM. In the case of PCC, the prediction accuracy decreases substantially when four or five features are used, compared to using just three features. This suggests that including features with lower importance ranks may lead to a decrease in accuracy due to increased complexity rather than a model improvement. Furthermore, when using SHAP, it can be explained that the Transformer model does not effectively capture the patterns required for time-series prediction.
[…]
Figure 7: The SOH prediction of applying the key feature set of B0005 identified by PCC to other cells ((a): B0006, (b): B0007, (c): B0018) and (d) the SOH prediction of applying the key feature set of B0006 identified by SHAP to B0005
4.3 Robust feature engineering to account for CtCV
In practical BMS applications in EVs industry, battery modules or packs consist of dozens or hundreds of cells. Therefore, performing feature engineering individually for all these cells is practically impossible. Selecting a common set of key features for cells with the same battery chemistry will significantly contribute to practical SOH prediction. Fig. 7(a-c) shows the results of applying the key features (i.e., F2, F10, F11) identified by PCC on the B0005 cell to SOH prediction for B0006, B0007, and B0018 cell. The prediction accuracy decreases significantly when compared to the accuracy achieved by training with the unique set of key features identified by applying PCC to each individual cell (Fig. 6(a-c)). On the other hand, when applying SHAP to identify the key features (i.e., F2, F11, F18) on B0006, B0007, and B0018 cell and then applying them to B0005 cell, the prediction accuracy is significantly higher (Fig. 7(d)). One possible interpretation of these results is that using the same key features can yield acceptable performance for cells with the same battery chemistry. This implies that feature selection using SHAP offers robust feature selection results that can account for CtCV.
@@ 4.3 Robust feature engineering to account for CtCV
Figure 8: Comparison of SOH predictions between feature selection methods according to training cycle: (a) 70 cycle, (b) 50 cycle, (c) 40 cycle, (d) 30 cycle.
Despite the high robustness of the SHAP-based approach, the rapid decrease in prediction accuracy with reduced training cycles suggests the need for a feature selection method that maximizes both robustness and performance. When training for 70 cycles, all the cases using all features, PCC-based selection cases, and SHAP-based selection cases have fairly high prediction accuracies. When considering all features, there is no significant difference between the real SOH and predicted SOH even as the training cycles are reduced to 50, 40, and 30. However, when the number of training cycles decreases, the accuracy rapidly declines with SHAP. However, with PCC, predicted SOH remain very close to the real SOH with 50 cycles for training and still effectively track the trend of the real SOH when training cycles are reduced to 40 or 30. This suggests that PCC might be more appropriate for predicting the SOH of individual cells and during extremely early cycles, while SHAP could be more suitable for predictions at the module or pack scale.
@@ 5 Conclusion
Accurate and efficient prediction of SOH is essential for realizing condition-based maintenance in EVs applications. In this work, we proposed an explainable SOH prediction framework combining feature engineering and the DLinear model. Leveraging public battery aging datasets, we demonstrated that DLinear outperforms conventional LSTM and Transformer models in both prediction accuracy and computational efficiency, even when trained with limited cycling data. To address the challenge of CtCV, we employed Pearson correlation and SHAP-based sensitivity analyses for feature selection. While PCC offered strong performance for individual cells, SHAP consistently identified key features applicable across multiple cells, enabling generalized SOH prediction for EVs battery modules. This is especially critical in large-scale battery systems, where individual cell-level modeling is impractical. The proposed framework offers a scalable, interpretable, and low-complexity solution with significant practical advantages for EVs BMS. The feature engineering process facilitates the generation of a dataset that is optimized for ML model training by enabling a substantial reduction in the number of features a BMS needs to measure and store. Furthermore, the use of SHAP ensures that key factors like CtCV are considered, while the DLinear model’s linear structure provides a substantially lower computational burden than more complex architectures like Transformer and LSTM. It enables early and reliable detection of battery degradation trends, contributing to safer operation, reduced maintenance costs, and improved battery utilization in electric vehicles. Future work will focus on integrating real-world EVs cycling data, refining model robustness under diverse usage conditions, and extending the framework for online learning in embedded BMS platforms.
@@
Appendix A Feature selection of B0005 cell
Figure A-1: Two temperature parameters (F10 and F13) of B0005 cell
Figure A-2: DLinear-based prediction results of B0005 cell using the two most sensitive parameters identified by SHAP